Briana and her sister Molly both wants to buy the same model bicycle. Briana needs $73 more before she can afford the bike. Molly needs $65 more. If they combine their money, they will have just enough to buy one bike that they could share. What is the cost of the bike?
Answer by richwmiller(17219)

b-73+b-65=b
2b-138=b
b-138=0
b=138